A proton projected in magnetic field of 0.020 T travels along a helical path of radius 5.0 cm and pitch 20 cm. Find the components of the velocity of the proton along and perpendicular to the magnetic field. Take the mass of the proton = 1.6 × 10^-27 kg.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. B ⃗ = 0.020 T MP= 1.6 × 10^–27 Kg Pitch = 20 cm = 2 × 10^–1 m Radius = 5 cm = 5 × 10^–2 m We know for a helical path, the velocity of the proton has got two components θ base ⊥ & θ base H Now, r = mθ⊥/qB ⇒ 5 * 10^-12 = 1.6 * 10^-27 * θ⊥/1.6 * 10^-19 *2 * 10^-2 ⇒ θ⊥ = 5 * 10^2 *1.6 * 10^-19 *2 * 10^-2/1.6 * 10^-27 = 1 * 10^5 m/s However, θ base H remains constant T = 2πm/qB Pitch = θ base H *T or, θ base H = Pitch/T θ base H = 2 * 10^-1/2 * 3.14 * 1.6 * 10^27 * 1.6 * 10^-19 *2 * 10^-2 = 0. 6369 * 10^5 = 6.4 * 10^4 m/s